# Diffusion approximation of the stationary distribution of a two-level single server queue

### v5

@@ 5.1 Process limit
$\rho_{11}p_{1,i}-p_{1,i+1}=\rho_{11}p_{1,i-1}-p_{1,i}+\rho_{21}p_{2,i-1},\qquad i\in[1,\ell_{0}-1].$
Summing up both sides of this equation for $i\in[1,j]$ for $j\in[1,\ell_{0}-1]$, we have
$\rho_{11}p_{1,j}-p_{1,j+1}=\rho_{21}\sum_{i=1}^{j}p_{2,i-1},\qquad j\in[0,\ell_{0}-1],$ (A.16)
because $\rho_{11}p_{1,0}-p_{1,1}=0$ by (A.1). This equation is valid also for $j=0$ by the convention that the empty sum vanishes, which we always take.
@@
$\rho_{11}p_{1,j}-p_{1,j+1}$ $=\rho_{21}\left(\rho_{21}^{-1}p_{2,1}+\sum_{i=1}^{j-1}\gamma_{12}^{\ell_{0}-i}p_{2,\ell_{0}}\right)$
$=\left(1+\rho_{21}\sum_{i=1}^{j-1}\gamma_{12}^{-(i-1)}\right)\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}}$
$=\left(1+\rho_{21}\frac{(\gamma_{12}^{-(j-1)}-1)\gamma_{12}}{1-\gamma_{12}}\right)\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}},\qquad j\in[1,\ell_{0}-1].$
Multiplying both sides by $\rho_{11}^{-(j+1)}$,
$\rho_{11}^{-j}p_{1,j}-\rho_{11}^{-(j+1)}p_{1,j+1}$
$\quad=\rho_{11}^{-(j+1)}\left(1+\rho_{21}\frac{(\gamma_{12}^{-(j-1)}-1)\gamma_{12}}{1-\gamma_{12}}1(j\geq 1)\right)\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}},\qquad j\in[0,\ell_{0}-1].$
Summing up both sides of this equation from $j=0$ to $j=k-1\leq\ell_{0}-1$, we have
$p_{1,0}-\rho_{11}^{-k}p_{1,k}=\sum_{j=0}^{k-1}\rho_{11}^{-(j+1)}\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}}+\sum_{j=1}^{k-1}\frac{(\rho_{11}\gamma_{12})^{-(j-1)}-\rho_{11}^{-(j-1)}}{1-\gamma_{12}}\rho_{11}^{-2}\rho_{21}\gamma_{12}^{\ell_{0}}p_{2,\ell_{0}}$
@@
We now apply these computations to the $r$-th 2-level $M/M/1$ queue, and compute the limit of its scaled stationary distribution through moment generating functions. For the $r$-th system, we put the superscript (r) as we have used for the $r$-th 2-level $GI/G/1$ queue. For example, $p^{(r)}_{m,i}$ is the stationary probability of $\{(J,L)=(m,i)\}$ for $(m,i)\in S$.
Define the moment generating functions $g^{(r)}_{m}(\theta)$ for $m=1,2,3$ as
$g^{(r)}_{m}(\theta)=\sum_{k=0}^{\ell^{(r)}_{0}}e^{r\theta k}p^{(r)}_{m,k},\quad\theta\in\mathbb{R},m=1,2,\qquad g^{(r)}_{3}(\theta)=\sum_{k=\ell^{(r)}_{0}+1}^{\infty}e^{r\theta k}p^{(r)}_{3,k},\quad\theta\leq 0,$
and define $g^{(r)}(\theta)=\sum_{m=1}^{3}g^{(r)}_{m}(\theta)$. Obviously, $g^{(r)}(\theta)$ is the moment generating function of $rL^{(r)}$. Thus, our task is to compute the limit of $g^{(r)}(\theta)$ as $r\downarrow 0$. To this end, we compute $g^{(r)}_{m}(\theta)$ for $m=1,2,3$, using the normalization condition that $g^{(r)}(0)=1$. However, it should be noted that this condition is not sufficient to prove the weak convergence of the distribution of $rL^{(r)}$, for which it is required that $\lim_{\theta\uparrow 0}\lim_{r\downarrow 0}g^{(r)}(\theta)=1$. This weak convergence will be verified through explicitly computing $g^{(r)}_{m}(\theta)$ for $m=1,2,3$.
We first compute $g^{(r)}_{m}(\theta)$ for $m=2,3$ because they are easier. It is immediate from (A.11), (A.10) and (A.9) that
$g_{2}^{(r)}(\theta)$ $=p^{(r)}_{2,0}+e^{\ell^{(r)}_{0}r\theta}\sum_{i=1}^{\ell^{(r)}_{0}}e^{-(\ell^{(r)}_{0}-i)r\theta}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}-i}p^{(r)}_{2,\ell^{(r)}_{0}}$
@@
We now compute the limits of $g^{(r)}_{m}(0)$ for $m=1,2,3$. For this, we note that the following asymptotic properties are obtained from the assumptions (3.b)–(3.e),
@@
$\lim_{r\downarrow 0}(1-\rho^{(r)}_{i})/r=\beta_{i},\qquad i=1,2,$ (A.23)
@@
where recall that $\beta_{i}=b_{i}/c_{i}$. Another important asymptotic properties are the ratios of the boundary probabilities. They are presented in the next lemma.

### v6

@@ 5.1 Process limit
$\rho_{11}p_{1,\ell}-p_{1,\ell+1}=\rho_{11}p_{1,\ell-1}-p_{1,\ell}+\rho_{21}p_{2,\ell-1},\qquad\ell\in[1,\ell_{0}-1].$
Summing up both sides of this equation for $\ell\in[1,m]$ for $m\in[1,\ell_{0}-1]$, we have
$\rho_{11}p_{1,m}-p_{1,m+1}=\rho_{21}\sum_{\ell=1}^{m}p_{2,\ell-1},\qquad m\in[0,\ell_{0}-1],$ (A.16)
because $\rho_{11}p_{1,0}-p_{1,1}=0$ by (A.1). This equation is valid also for $m=0$ by the convention that the empty sum vanishes, which we always take.
@@
$\rho_{11}p_{1,m}-p_{1,m+1}$ $=\rho_{21}\left(\rho_{21}^{-1}p_{2,1}+\sum_{\ell=1}^{m-1}\gamma_{12}^{\ell_{0}-\ell}p_{2,\ell_{0}}\right)$
$=\left(1+\rho_{21}\sum_{\ell=1}^{m-1}\gamma_{12}^{-(\ell-1)}\right)\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}}$
$=\left(1+\rho_{21}\frac{(\gamma_{12}^{-(m-1)}-1)\gamma_{12}}{1-\gamma_{12}}\right)\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}},\qquad m\in[1,\ell_{0}-1].$
Multiplying both sides by $\rho_{11}^{-(m+1)}$,
$\rho_{11}^{-m}p_{1,m}-\rho_{11}^{-(m+1)}p_{1,m+1}$
$\quad=\rho_{11}^{-(m+1)}\left(1+\rho_{21}\frac{(\gamma_{12}^{-(m-1)}-1)\gamma_{12}}{1-\gamma_{12}}1(m\geq 1)\right)\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}},\qquad m\in[0,\ell_{0}-1].$
Summing up both sides of this equation from $m=0$ to $m=k-1\leq\ell_{0}-1$, we have
$p_{1,0}-\rho_{11}^{-k}p_{1,k}=\sum_{m=0}^{k-1}\rho_{11}^{-(m+1)}\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}}+\sum_{m=1}^{k-1}\frac{(\rho_{11}\gamma_{12})^{-(m-1)}-\rho_{11}^{-(m-1)}}{1-\gamma_{12}}\rho_{11}^{-2}\rho_{21}\gamma_{12}^{\ell_{0}}p_{2,\ell_{0}}$
@@
We now apply these computations to the $r$-th 2-level $M/M/1$ queue, and compute the limit of its scaled stationary distribution through moment generating functions. For the $r$-th system, we put the superscript (r) as we have used for the $r$-th 2-level $GI/G/1$ queue. For example, $p^{(r)}_{j,\ell}$ is the stationary probability of $\{(J,L)=(j,\ell)\}$ for $(j,\ell)\in S$.
Define the moment generating functions $g^{(r)}_{j}(\theta)$ for $j=1,2,3$ as
$g^{(r)}_{j}(\theta)=\sum_{k=0}^{\ell^{(r)}_{0}}e^{r\theta k}p^{(r)}_{j,k},\quad\theta\in\mathbb{R},j=1,2,\qquad g^{(r)}_{3}(\theta)=\sum_{k=\ell^{(r)}_{0}+1}^{\infty}e^{r\theta k}p^{(r)}_{3,k},\quad\theta\leq 0,$
and define $g^{(r)}(\theta)=\sum_{j=1}^{3}g^{(r)}_{j}(\theta)$. Obviously, $g^{(r)}(\theta)$ is the moment generating function of $rL^{(r)}$. Thus, our task is to compute the limit of $g^{(r)}(\theta)$ as $r\downarrow 0$. To this end, we compute $g^{(r)}_{j}(\theta)$ for $j=1,2,3$, using the normalization condition that $g^{(r)}(0)=1$. However, it should be noted that this condition is not sufficient to prove the weak convergence of the distribution of $rL^{(r)}$, for which it is required that $\lim_{\theta\uparrow 0}\lim_{r\downarrow 0}g^{(r)}(\theta)=1$. This convergence will be verified through explicitly computing $g^{(r)}_{j}(\theta)$ for $j=1,2,3$.
We first compute $g^{(r)}_{j}(\theta)$ for $j=2,3$ because they are easier. It is immediate from (A.11), (A.10) and (A.9) that
$g_{2}^{(r)}(\theta)$ $=p^{(r)}_{2,0}+e^{\ell^{(r)}_{0}r\theta}\sum_{\ell=1}^{\ell^{(r)}_{0}}e^{-(\ell^{(r)}_{0}-\ell)r\theta}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}-\ell}p^{(r)}_{2,\ell^{(r)}_{0}}$
@@
We now compute the limits of $g^{(r)}_{j}(0)$ for $j=1,2,3$. For this, we note that the following asymptotic properties are obtained from the assumptions (3.b)–(3.e),
@@
$\lim_{r\downarrow 0}(1-\rho^{(r)}_{\ell})/r=\beta_{\ell},\qquad\ell=1,2,$ (A.23)
@@
where recall that $\beta_{\ell}=b_{\ell}/c_{\ell}$. Another important asymptotic properties are the ratios of the boundary probabilities. They are presented in the next lemma.